OpenStudy (anonymous):

Express cos x in terms of tan x/2

OpenStudy (anonymous):

I suppose you could use the tangent half-angle identity plus some other identity to do that..

OpenStudy (anonymous):

From \[\tan(x/2)=\frac{\sin x}{1+\cos x} \rightarrow \cos x =\frac{\sin x}{\tan(x/2)} \space -1\]

OpenStudy (anonymous):

Then you could use another identity to get sin x out of there.

OpenStudy (anonymous):

Thank you so much for getting me started. I really appreciate the help.

OpenStudy (anonymous):

Oh wait, you completly answered my question. Cos x is expressed in tan x/2. Again, I really appreciate the help.

OpenStudy (anonymous):

Hmm, maybe. I wouldn't be satisfied unless the only function on the right side was tan(x/2). It is pretty tricky to get that sin x out of there though, so if you're satisfied, I'm satisfied!